If 300 jellybeans cost you x dollars. How many jellybeans can you purchase for 50 cents at the same rate?
I know the answer is 150/x, but I can't understand why?!
Thank you

Answer by CharlesG2(834)

If 300 jellybeans cost you x dollars. How many jellybeans can you purchase for 50 cents at the same rate?
I know the answer is 150/x, but I can't understand why?!
Thank you
300 JB / x dollars = y JB / 0.5 dollars
300 * 0.5 = x * y
150/x = y